There are two naturally occurring isotopes of copper. 63Cu has a mass of 62.9296 amu. 65Cu has a mass of 64.9278 amu. Determine the abundance of each isotope. Number 63 0 Number 65 0
Answer #1

MW of copper = 63.546 g/mol

MW average = MW of copper = 63.546 g/mol

also

MW avg = x1*MW1 + x"*MW2

63.546 = x1*(62.9296) + (1-x1)*64.9278

63.546 = 62.9296*x1 + 64.9278 - 64.9278x1

63.546 -64.9278 = (62.9296-64.9278) * x1

(63.546 -64.9278 ) / (62.9296-64.9278) = x1

x1 = 0.69152

x2 = 1-x1 = 1-0.69152 = 0.30848

then

63 isotope --> 69.15%

63 isotope --> 30.85%